OpenStudy (anonymous):

what method is easiest when it comes to 2x^2 - x - 6?

OpenStudy (anonymous):

factoring...

OpenStudy (anonymous):

factoring

OpenStudy (anonymous):

what factoring method helps

OpenStudy (anonymous):

(x-2)(2x+3) x= 2 x=-3/2

OpenStudy (anonymous):

so u just add 2x instead of just x?

OpenStudy (anonymous):

2x*x = 2x^2 right ?

OpenStudy (anonymous):

-6= -2*3 right?

OpenStudy (anonymous):

ya